OpenStudy (anonymous):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 2, – 16), and x intercepts at x = -6 and x = 2. y = x2 + 4x − ___

OpenStudy (anonymous):

You can plug in your vertex (-2,-16) for x and y and solve for the unknown number

OpenStudy (anonymous):

so it will be -68?

OpenStudy (anonymous):

No. I'll write it out

OpenStudy (anonymous):

\[y=x^2+4x-A\] A is the unknown. Plug in y=-16 and x=-2

OpenStudy (anonymous):

Let me know if u need it further

OpenStudy (anonymous):

y= -60 - A??

OpenStudy (anonymous):

Plug in y=-16 and x=-2 \[-16=(-2)^2+4(-2)-A\] -16=4-8-A Add 'A' to both sides -16+A=-4-A+A 'A' cancels on the right, so u have A-16=-4 Add 16 to both sides A-16+16=-4+16 16 npw cancels on the left, so u have A=-4+16 A=12

OpenStudy (anonymous):

so the answer is 12?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank youu

OpenStudy (anonymous):

LOL Your welcome. You should try and write out the steps next time

OpenStudy (anonymous):

alright ill remember too